OpenStudy (anonymous):

6r + r - 5r A. 2r B. 1r + r C. 0r D. 7r - 5r

OpenStudy (anonymous):

C

OpenStudy (agent0smith):

Would you know the answer if it was 6 + 1 - 5 ?

OpenStudy (nincompoop):

all r's can be added and subtracted together

OpenStudy (anonymous):

The simplified expression

OpenStudy (agent0smith):

6 + 1 - 5 =???

OpenStudy (nincompoop):

pretend r is the rubber one must use to make a cork how many rubbers are there?

OpenStudy (anonymous):

Its B.

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

i got 3 different answers, so can someone prove their point.

OpenStudy (agent0smith):

Answer this 6 + 1 - 5 =

OpenStudy (anonymous):

2

OpenStudy (agent0smith):

Good, so 6r + 1r - 5r =

OpenStudy (agent0smith):

The r is really 1r.

OpenStudy (anonymous):

You work from left to right so it would be 7r-5r

OpenStudy (anonymous):

You can't add r to 6r or 5r. You would subtract 5r from 6r ( both are like variables) and get 1r. 6r-5r+r =1r+r

OpenStudy (agent0smith):

@Unicornkat and @Baseballa101 stop it.

OpenStudy (anonymous):

Oh so it's B?

OpenStudy (agent0smith):

6 + 1 - 5 = 2 so 6r + 1r - 5r =

OpenStudy (anonymous):

yes. Its B

OpenStudy (anonymous):

WHAT LETTER IS IT AGENT SMITH.

OpenStudy (anonymous):

;-;

OpenStudy (agent0smith):

Look at the above post, NO IT IS NOT B.

OpenStudy (anonymous):

It's A, m'dear

OpenStudy (agent0smith):

6 + 1 - 5 = 2 so 6r + 1r - 5r = Here is an example: 5m + 4m - 3m = 6m since 5+4-3 is 6.

OpenStudy (anonymous):

6r+1r = 7r 7r-5r = 2r It's A

OpenStudy (agent0smith):

btw @UnicornKat "You can't add r to 6r or 5r." Yes you can. r is 1r, which can be added to 6r or 5r. They're all like terms.

OpenStudy (anonymous):

Oh whatever @agent0smith

OpenStudy (agent0smith):

lol that's a good response, quite helpful.